# Investigating FRB 20240114A with FAST: Morphological Classification and Drifting Rate Measurements in a Burst-Cluster Framework

### v1

@@ 2.1 Data Analysis and Burst Detection
Furthermore, to improve data quality and effectively remove RFI, we enabled the --zdot and --kadane options during the replot_fil process. The --zdot option was used to remove RFI with a DM of 0, typically caused by terrestrial radio sources and unrelated to genuine astrophysical signals. The --kadane option applied the Kadane algorithm, which uses a maximum subarray summation method to detect and remove high-power anomalous signals. This method efficiently identifies and removes RFI, minimizing the impact of spurious signals and improving the quality of the data for subsequent analysis. Details about the algorithms can be found in Men et al. (2019, 2023).
2.2 DM Optimized for Structure
The radio signal experiences dispersion as it propagates through the plasma. This dispersive delay is inversely proportional to the square of the radio frequency and directly proportional to the DM, which quantifies the total plasma content along the signal’s path:
@@ 2.2 DM Optimized for Structure
where $\mathrm{DM}$ is the dispersion measure, $n$ is the free electron density, and $dl$ is the differential path length.
The FRB search algorithm explores a range of DM values to compensate for the delay. According to Hessels et al. (2019), the DM of the FRB is determined based on either the maximum signal-to-noise ratio (S/N) or the most detailed temporal structure of the burst. Motivated by the propagation model proposed by Tuntsov et al. (2021), the intrinsic burst emission passes through an inhomogeneous modulation region near the FRB source, where broadband but temporally sharp modulation features are imprinted onto the already dispersed pulse. Assuming no significant source-inherent drifting features exist, the modulated temporal structure imprinted by the screen can be optimally resolved by adopting a structure-maximizing DM, whereas the maximum S/N-based DM predominantly captures dispersion accumulated along the entire source-Earth path. Conversely, when propagation effects are negligible, the de-dispersed dynamic spectrum morphology directly manifests the source’s intrinsic radiation mechanism. The interpretation becomes more nuanced when both propagation-induced modulation and source-inherent drifting features coexist.
[…]
Figure 2: Both figures compare the DM_power results for two data sets: March 11, 2024 session (20240311) and March 13, 2024 session (20240313), obtained using the DM_power software. The average DM values for these two days are shown for comparison. Interestingly, the DM value for the March 12, 2024 (20240312) data lies precisely between the values of the March 11 and March 13 datasets, suggesting that the DM for this day is consistent with a midpoint between the two neighboring observations.
3 burst-cluster Morphological Classification
Table 1: The morphological classification results after de-dispersion using the average DM value (529.2 pc cm-3). U and D denote upward and downward, respectively. H, M, L, and W represent high, middle, low, and wide frequency bands (based on the 1-1.5 GHz observation range of the FAST telescope). The numbers 1, 2, and m indicate single-, double-, and multiple-components burst-clusters, respectively. C denotes complex morphology, NE represents burst-clusters with no evidence of drifting, and Not-Clear indicates cases with low signal-to-noise ratio or severe RFI, making it impossible to determine their morphology. ND stands for no drifting.
@@ C NE-H NE-L Not-Clear No drifting (ND=720)
In this section, we detail our morphological classification method for bursts, which adapts the framework of Zhou et al. (2022) with modifications. However, it should be noted that there is a major caveat regarding our DM averaging approach, which may affect the identification of drifting burst-clusters, particularly for single-component burst-clusters where the observed drifting may be an artifact rather than genuine frequency evolution. A follow-up study using precise DM values for all FAST observations of FRB 20240114A is currently underway to address this limitation. Our morphological classification criteria are defined as follows (for a more detailed schematic illustration, see Figure 3):
@@
    Burst: A group of sub-bursts, where the emission intensity between adjacent sub-bursts remains above the background level and exhibits a distinguished peak (visually separable from background fluctuations) in the de-dispersed burst profile. In particular, a burst consisting of only a sub-burst is also considered a burst.
@@
    Burst-Cluster: A collection of bursts (excluding those labeled as Not-Clear), with adjacent components (or peaks) separations is less than the time threshold $\tau$, which is the valley derived from the bimodal waiting-time distribution. In particular, a single burst is still classified as a burst-cluster if no other bursts are detected within the $\tau$-duration intervals immediately preceding and following it. Here, the time threshold $\tau$ is 400 ms for FRB 20240114A, as derived from the bimodal waiting-time distribution shown in Jun-Shuo Zhang (2025) (see their Extended Data Figure 2 for the distribution). Each entry also excluding Not-Clear events in Table 1 represents one such burst-cluster.
The counts for single-, double-, and multiple-components burst-clusters (1, 2, and m) in Table 1 only consider the number of sub-bursts of same burst-cluster. In contrast, the 1, 2, and m counts defined by Zhou et al. (2022) specifically refer to single-, double-, and multiple-component bursts (rather than burst-clusters), focusing on the drifting of sub-bursts within the same burst. Therefore, a single-component burst-cluster contains only one burst, and this burst only contains one sub-burst. A double-components burst-cluster may consist of either one burst containing two sub-bursts, or two bursts each containing one sub-burst. By extension, a multi-components burst-cluster can manifest as either one burst with multiple sub-bursts, or multiple bursts containing varied numbers of sub-bursts.
Table 1 displays the de-dispersed morphological classifications that were derived by applying the mean DM of 529.2 pc cm-3. In the table, D represents the downward drifting and U represents the upward drifting. H, M, L, and W denote high-frequency band, middle-frequency band, low-frequency band, and wide-frequency band of at least 500 MHz, respectively. These frequency bands are defined based on the observation range of the FAST (1.0-1.5 GHz). The numbers 1, 2, and m represent single-, double-, and multiple-components burst-cluster, respectively. C indicates complex morphology, NE-H stands for no-evidence high frequency band, NE-L stands for no-evidence low frequency band, and Not-Clear indicates cases where the signal-to-noise ratio is too low or the radio frequency interference (RFI) is too severe. ND represents no drifting.
@@
According to the definitions described above, as shown in Figure 3, burst-cluster A consists of three bursts (including two intermittent time intervals). The first burst contains three sub-bursts (including two consecutive time intervals), the second burst contains only one sub-burst (with no consecutive time intervals), and the third burst contains two sub-bursts (including one consecutive time interval). In total, this results in six sub-bursts, three consecutive time intervals, and two intermittent time intervals. Similarly, burst-cluster B contains one burst, one sub-burst, zero consecutive time intervals, and zero intermittent time intervals. Burst-cluster C contains five bursts, six sub-bursts, one consecutive time interval, and four intermittent time intervals. Notably, in burst-cluster C, although the first two sub-bursts are closely connected, the emission intensity between them returns to the background level; thus according to our definition, these two sub-bursts belong to different bursts. Burst-cluster D contains one burst, two sub-bursts, one consecutive time interval, and zero intermittent time intervals. Therefore, in this observation session, the overall counts of burst-clusters, bursts, sub‐bursts, consecutive time intervals, and intermittent time intervals are four, ten, fifteen, five and six, respectively. For a more comprehensive analysis, the distribution of consecutive and intermittent time intervals for the 978 burst-clusters exhibiting drifting is presented in Section 4.3.
3.1 Downward Drifting burst-clusters
@@
Downward drifting burst-clusters refer to cases where the higher-frequency components of the burst-cluster arrive earlier than the lower-frequency ones. These burst-clusters are categorized based on the number of components (sub-bursts) observed in their dynamic spectrum and the position of the frequency bands: high-frequency band (H), middle-frequency band (M), low-frequency band (L), and wide-frequency band (W), as recorded by the FAST. For single-component burst-clusters, this implies that the higher-frequency sub-band arrives earlier than the lower-frequency one. In the case of double-component burst-clusters, the sub-burst with the higher peak frequency is detected earlier than the one with the lower peak frequency. For multiple burst-clusters, the sub-burst with the highest peak frequency is the first to arrive, while the sub-burst with the lowest peak frequency arrives last, which facilitates the calculation of the drifting rate. A representation of these burst-clusters can be found in Figure 4.
@@
As shown in Table 1, a total of 745 downward drifting burst-clusters were detected. Among them, there were 364 single-component burst-clusters (48.9%). It should be noted that since we used an average DM value rather than the precise DM for each individual burst-cluster, the apparent drifting in single-component burst-clusters could be an artifact of this DM averaging approach, and the true number of genuinely drifting single-component burst-clusters could be much smaller than what we report here. Additionally, there were 334 double-components burst-clusters (44.8%), and 47 multiple-components burst-clusters (6.3%). The distribution across frequency bands is as follows: 181 burst-clusters (24.3%) in the high-frequency band (H), 60 burst-clusters (8.1%) in the middle-frequency band (M), 300 burst-clusters (40.3%) in the low-frequency band (L), and 204 burst-clusters (27.4%) in the wide-frequency band (W). The number of consecutive time intervals for downward drifting burst-clusters was 184, and the number of intermittent time intervals for downward drifting burst-clusters was 265.
@@
The prevalent downward drifting can be explained by multiple physical frameworks. The most natural interpretation invokes radius to frequency mapping in magnetospheric models (Wang et al., 2019; Lyutikov, 2020; Zhang, 2022). The curvature radiation model suggests that relativistic particle bunches moving along neutron star open field lines naturally exhibit downward drifting, as emission regions at higher altitudes encounter larger curvature radii (Wang et al., 2019, 2020). Separately, relativistic kinematics introduces a distinct mechanism: bulk motion combined with Dicke’s superradiance predicts an inverse correlation between sub-burst drifting rates and durations through Doppler-shifted narrowband emission (Rajabi et al., 2020). Propagation effects also contribute significantly; chromatic deflection by circum-burst plasma prisms (Tuntsov et al., 2021) can redirect low-frequency components into the observer’s line of sight, mimicking downward drifting. Synchrotron maser emission from decelerating shocks (Metzger et al., 2019, 2022) provides an alternative mechanism where shock cooling drives $\nu_{c}\propto t^{-\beta}$ decay, consistent with the majority of drifting events.
3.2 Upward Drifting burst-clusters
@@
Figure 5: The typical dynamic spectra of upward drifting burst-clusters. Each burst-cluster is labeled as Ux-y, where x denotes the number of components (labeled as ”m” if the number exceeds two), and y indicates the observation band: L (low frequency), M (middle frequency), H (high frequency), W (wide frequency). The plots are arranged in a grid of four rows and three columns, with x corresponding to each column and y corresponding to each row. Notably, our sample contains no events classified as Um-M or U1-W.
@@
Upward drifting burst-clusters refer to cases where the lower-frequency components of the burst-cluster precede the higher-frequency components. These burst-clusters are classified based on the number of components (or sub-bursts) observed in their dynamic spectrum and the position of the frequency bands: high-frequency band (H), middle-frequency band (M), low-frequency band (L), and wide-frequency band (W), as recorded by the FAST. For single-component burst-clusters, this means that the lower-frequency sub-band arrives earlier than the higher-frequency one. In the case of double-components burst-clusters, the sub-burst with the lower peak frequency is detected earlier than the one with the higher peak frequency. For multiple-components burst-clusters, the sub-burst with the lowest peak frequency arrives first, while the sub-burst with the highest peak frequency arrives last, allowing for the measurement of the drifting rate. It should be noted that some categories, such as U1-W and Um-M, currently lack samples due to insufficient data. An illustration of these burst-clusters can be seen in Figure 5.
@@
As shown in Table 1, a total of 233 upward drifting burst-clusters were detected. Among them, there were 142 single-component burst-clusters (60.9%), with the same caveat regarding DM averaging effects as mentioned above for downward drifting cases. Additionally, there were 86 double-components burst-clusters (36.9%), and 5 multiple-components burst-clusters (2.1%). The distribution across frequency bands is as follows: 45 burst-clusters (19.3%) in the high-frequency band (H), 47 burst-clusters (20.2%) in the middle-frequency band (M), 117 burst-clusters (50.2%) in the low-frequency band (L), and 27 burst-clusters (11.6%) in the wide-frequency band (W). The number of consecutive time intervals for upward drifting burst-clusters was 90, and the number of intermittent time intervals for upward drifting burst-clusters was 9.

### v3

@@ II.1 Data Analysis and Burst Detection
Furthermore, to mitigate RFI and obtain cleaner data for subsequent analysis, we enabled the --zdot and --kadane options during the replot_fil process. The --zdot option was used to remove zero-DM RFI, typically caused by terrestrial radio sources and unrelated to genuine astrophysical signals. The --kadane option applied the Kadane algorithm, which uses a maximum subarray summation method to detect and remove high-power anomalous signals. This method efficiently identifies and removes RFI, minimizing the impact of spurious signals for subsequent analysis. Details about the algorithms can be found in 2019MNRAS.488.3957M; 2023A&A...679A..20M.
II.2 DM Optimized for Structure
The radio signal experiences dispersion as it propagates through the plasma. This dispersive delay is inversely proportional to the square of the radio frequency and directly proportional to the dispersion measure (DM), which quantifies the total plasma content along the line of sight from the source to the observer:
$\mathrm{DM}=\int n{\rm d}l.$ (2)
where $n$ is the free electron density, and $dl$ is the differential path length.
The FRB search algorithm explores a range of DM values to compensate for the delay. According to Hessels_2019, the DM of the FRB is determined based on either the maximum signal-to-noise ratio (S/N) or the most detailed temporal structure of the burst. As described by the propagation scenario of 2021arXiv210713549T, the intrinsic burst emission passes through an inhomogeneous modulation region near the FRB source, where short-duration intensity dropouts or narrow absorption features are imprinted simultaneously across the full observing band onto the already dispersed pulse. Assuming no significant source-inherent drifting features exist, the modulated temporal structure imprinted by the screen can be optimally resolved by adopting a structure-maximizing DM, whereas the S/N-maximizing DM corresponds to the value that yields the highest integrated burst significance, though this choice may smear out complex temporal substructures or multiple components. Conversely, when propagation effects are negligible, the de-dispersed dynamic spectrum morphology directly manifests the source’s intrinsic radiation mechanism. The interpretation becomes more nuanced when both propagation-induced modulation and source-inherent drifting features coexist.
@@ II.2 DM Optimized for Structure
Figure 1: Comparison of two DM measurement methods, DM_phase and DM_power, applied to the data from March 12, 2024 (20240312). The left panel shows results from DM_phase, while the right panel corresponds to DM_power, both yielding consistent results. The dotted line indicates the average DM value, denoted by ⟨DM⟩, which is determined to be $529.2~\text{pc~cm}^{-3}$ for both DM_phase and DM_power, with the standard error (SE) reflecting the variability in the measurements. Only data points with error bars $1~\text{pc~cm}^{-3}$ (i.e., a one-sided error bar of $0.5~\text{pc~cm}^{-3}$) are plotted, selected from the original 3203 bursts, with the number of precise data points noted in Sample Count.
Figure 2: Both figures compare the DM_power results for two data sets: March 11 2024 session (20240311) and March 13 2024 session (20240313), obtained using the DM_power software. The average DM values for these two days are shown for comparison. Interestingly, the DM value for the March 12, 2024 (20240312) data lies precisely between the values of the March 11 and March 13 datasets, suggesting that the DM for this day is consistent with a midpoint between the two neighboring observations.
III burst-cluster Morphological Classification
Table 1: The morphological classification results after de-dispersion using the average DM value (529.2 pc cm-3). U and D denote upward and downward, respectively. H, M, L, and W represent high, middle, low, and wide frequency bands (based on the 1-1.5 GHz observation range of the FAST telescope). The numbers 1, 2, and m indicate single-, double-, and multi-component burst-clusters, respectively. C denotes complex morphology, NE represents burst-clusters with no evidence of drifting, and Not-Clear indicates cases with low signal-to-noise ratio or severe RFI, making it impossible to determine their morphology. ND stands for no drifting.
@@ C NE-H NE-L Not-Clear No drifting (ND=720)
In this section, we present our morphological classification method for bursts, which is adapted from the framework of 2022RAA....22l4001Z with modifications. However, it should be noted that there is a major caveat regarding our DM averaging approach, which may affect the identification of drifting burst-clusters, particularly for single-component burst-clusters where the observed drifting may be an artifact rather than genuine frequency evolution. A follow-up study using precise DM values for all FAST observations of FRB 20240114A is currently underway to address this limitation. Our morphological classification criteria are defined as follows (for a more detailed schematic illustration, see Figure 3):
@@
    Burst: A collection of sub-bursts, which may consist of only one sub-burst.
@@
    Burst-Cluster: A collection of bursts (excluding those labeled as Not-Clear), with adjacent components (or peaks) separations less than the time threshold $\tau$, which is the valley derived from the bimodal waiting-time distribution. In particular, a single burst is still classified as a burst-cluster if no other bursts are detected within the $\tau$-duration intervals immediately preceding and following it. Here, the time threshold $\tau$ is 400 ms for FRB 20240114A, as derived from the bimodal waiting-time distribution shown in 2025arXiv250714707Z (see their Extended Data Fig. 2 for the distribution). Each entry excluding Not-Clear events in Table 1 represents one such burst-cluster.
The counts for single-, double-, and multi-component burst-clusters (1, 2, and m) in Table 1 only consider the number of sub-bursts of same burst-cluster. In contrast, the 1, 2, and m counts defined by 2022RAA....22l4001Z specifically refer to single-, double-, and multi-component bursts (rather than burst-clusters), focusing on the drifting of sub-bursts within the same burst. Therefore, a single-component burst-cluster contains only one burst, and this burst only contains one sub-burst. A double-component burst-cluster may consist of either one burst containing two sub-bursts, or two bursts each containing one sub-burst. By extension, a multi-component burst-cluster can manifest as either one burst with multiple sub-bursts, or multiple bursts containing varied numbers of sub-bursts.
Table 1 displays the de-dispersed morphological classifications that were derived by applying the mean DM of 529.2 pc cm-3. In the table, D represents the downward drifting and U represents the upward drifting. The codes H, M, L, and W classify emissions with a bandwidth relative to the FAST observing band (1.0-1.5 GHz). M (Middle): Emission is within the 1.0-1.5 GHz band. H (High): Emission extends above the 1.5 GHz upper boundary. L (Low): Emission extends below the 1.0 GHz lower boundary. W (Wide): Emission extends beyond both the 1.0 GHz and 1.5 GHz boundaries. The numbers 1, 2, and m represent single-, double-, and multi-component burst-cluster, respectively. C indicates complex morphology, NE-H stands for no-evidence high frequency band, NE-L stands for no-evidence low frequency band, and Not-Clear indicates cases where the signal-to-noise ratio is too low or the radio frequency interference (RFI) is too severe. ND represents no drifting.
@@
According to the definitions described above, as shown in Figure 3, burst-cluster A consists of three bursts (including two intermittent time intervals). The first burst contains three sub-bursts (including two consecutive time intervals), the second burst contains only one sub-burst (with no consecutive time intervals), and the third burst contains two sub-bursts (including one consecutive time interval). In total, this results in six sub-bursts, three consecutive time intervals, and two intermittent time intervals. Similarly, burst-cluster B contains one burst, one sub-burst, zero consecutive time intervals, and zero intermittent time intervals. Burst-cluster C contains five bursts, six sub-bursts, one consecutive time interval, and four intermittent time intervals. Notably, in burst-cluster C, although the first two sub-bursts are closely connected, the emission intensity between them returns to the background level; thus according to our definition, these two sub-bursts belong to different bursts. Burst-cluster D contains one burst, two sub-bursts, one consecutive time interval, and zero intermittent time intervals. Therefore, in this observation session, the overall counts of burst-clusters, bursts, sub‐bursts, consecutive time intervals, and intermittent time intervals are four, ten, fifteen, five and six, respectively. For a more comprehensive analysis, the distribution of consecutive and intermittent time intervals for the 978 burst-clusters exhibiting drifting is presented in Section IV.3.
III.1 Downward Drifting burst-clusters
@@
Downward drifting burst-clusters refer to cases where the higher-frequency components of the burst-cluster arrive earlier than the lower-frequency ones. These burst-clusters are categorized based on the number of components (sub-bursts) observed in their dynamic spectrum and the position of the frequency bands: high-frequency band (H), middle-frequency band (M), low-frequency band (L), and wide-frequency band (W), as recorded by the FAST. For single-component burst-clusters, this implies that the higher-frequency sub-band arrives earlier than the lower-frequency one. In the case of double-component burst-clusters, the sub-burst with the higher peak frequency is detected earlier than the one with the lower peak frequency. For multi-component burst-clusters, the sub-burst with the highest peak frequency is the first to arrive, while the sub-burst with the lowest peak frequency arrives last, which facilitates the calculation of the drifting rate. A representation of these burst-clusters can be found in Figure 4.
@@
As shown in Table 1, a total of 745 downward drifting burst-clusters were detected. Among them, there were 364 single-component burst-clusters (48.9%). It should be noted that since we used an average DM value rather than the precise DM for each individual burst-cluster, the apparent drifting in single-component burst-clusters could be an artifact of this DM averaging approach, and the true number of genuinely drifting single-component burst-clusters could be much smaller than what we report here. Additionally, there were 334 double-component burst-clusters (44.8%), and 47 multi-component burst-clusters (6.3%). The distribution across frequency bands is as follows: 181 burst-clusters (24.3%) in the high-frequency band (H), 60 burst-clusters (8.1%) in the middle-frequency band (M), 300 burst-clusters (40.3%) in the low-frequency band (L), and 204 burst-clusters (27.4%) in the wide-frequency band (W). The number of consecutive time intervals for downward drifting burst-clusters was 184, and the number of intermittent time intervals for downward drifting burst-clusters was 265.
@@
The prevalent downward drifting can be explained by multiple physical frameworks. The most natural interpretation invokes radius to frequency mapping in magnetospheric models (2019ApJ...876L..15W; osti_1802960; 2022ApJ...925...53Z). The curvature radiation model suggests that relativistic particle bunches moving along neutron star open field lines naturally exhibit downward drifting, as emission regions at higher altitudes encounter larger curvature radii (2019ApJ...876L..15W; 2020ApJ...899..109W). Separately, relativistic kinematics introduces a distinct mechanism: bulk motion combined with Dicke’s superradiance predicts an inverse correlation between sub-burst drifting rates and durations through Doppler-shifted narrowband emission (2020MNRAS.498.4936R). Propagation effects also contribute significantly; chromatic deflection by circum-burst plasma prisms (2021arXiv210713549T) can redirect low-frequency components into the observer’s line of sight, mimicking downward drifting. Synchrotron maser emission from decelerating shocks (10.1093/mnras/stz700; 2022ApJ...925..135M) provides an alternative mechanism where shock cooling drives $\nu_{c}\propto t^{-\beta}$ decay, consistent with the majority of drifting events. Taken together, these scenarios indicate that downward drifting is a natural and robust outcome of several distinct physical processes, though distinguishing their relative contributions remains challenging.
III.2 Upward Drifting burst-clusters
@@
Figure 5: The typical dynamic spectra of upward drifting burst-clusters. Each burst-cluster is labeled as Ux-y, where x denotes the number of components (labeled as ’m’ if the number exceeds two), and y indicates the observation band: L (low frequency), M (middle frequency), H (high frequency), W (wide frequency). The plots are arranged in a grid of four rows and three columns, with x corresponding to each column and y corresponding to each row. Notably, our sample contains no events classified as Um-M or U1-W.
@@
Upward drifting burst-clusters refer to cases where the lower-frequency components of the burst-cluster precede the higher-frequency components. These burst-clusters are classified based on the number of components (or sub-bursts) observed in their dynamic spectrum and the position of the frequency bands: high-frequency band (H), middle-frequency band (M), low-frequency band (L), and wide-frequency band (W), as recorded by the FAST. For single-component burst-clusters, this means that the lower-frequency sub-band arrives earlier than the higher-frequency one. In the case of double-component burst-clusters, the sub-burst with the lower peak frequency is detected earlier than the one with the higher peak frequency. For multi-component burst-clusters, the sub-burst with the lowest peak frequency arrives first, while the sub-burst with the highest peak frequency arrives last, allowing for the measurement of the drifting rate. It should be noted that some categories, such as U1-W and Um-M, currently lack samples due to insufficient data. An illustration of these burst-clusters can be seen in Figure 5.
@@
As shown in Table 1, a total of 233 upward drifting burst-clusters were detected. Among them, there were 142 single-component burst-clusters (60.9%), with the same caveat regarding DM averaging effects as mentioned above for downward drifting cases. Additionally, there were 86 double-component burst-clusters (36.9%), and 5 multi-component burst-clusters (2.1%). The distribution across frequency bands is as follows: 45 burst-clusters (19.3%) in the high-frequency band (H), 47 burst-clusters (20.2%) in the middle-frequency band (M), 117 burst-clusters (50.2%) in the low-frequency band (L), and 27 burst-clusters (11.6%) in the wide-frequency band (W). The number of consecutive time intervals for upward drifting burst-clusters was 90, and the number of intermittent time intervals for upward drifting burst-clusters was 9.